I am 4 times as old as my daughter, I was 39 when my daughter was born, How old am I now ?
Answer by longjonsilver(2297)

First, define...
Let x = my age now
Ley y = my daughter's age now.
so, x=4y ---eqn1
and x-y=39 -eqn2
so, sub 1 into 2, giving 4y-y=39
3y=39
y=13.
Hence x=4*13
x=52